# Landing with the Score: Riemannian Optimization through Denoising

### v1

@@ 7 Conclusion and Future Work
Appendix A More on manifolds and distance functions
@@
In addition to the notation introducted in Section 2.1, we note that $\tau_{\mathcal{M}}$ is also the largest $\tau\geq 0$ such that the map $\{(p,v)\in\operatorname{N}\mathcal{M}\mid\lVert v\rVert<\tau\}\to\mathbb{R}^{d}:(p,v)\mapsto p+v$ is a diffeomorphism. By a tubular neighborhood of radius $\tau\in(0,\tau_{\mathcal{M}}]$ we mean a set of the form $\mathcal{T}(\tau)=\{p+v\mid p\in\mathcal{M}\,,v\in\operatorname{N}_{p}\mathcal{M}\,,\lVert v\rVert<\tau\}$. Moreover, for $x\in\mathbb{R}^{d}$ let $\operatorname{dist}_{\mathcal{M}}(x)=\inf_{p\in\mathcal{M}}\lVert x-p\rVert$ denote the distance function so that $\operatorname{d}(x)=\frac{1}{2}\operatorname{dist}_{\mathcal{M}}(x)^{2}$. The second fundamental form of $\mathcal{M}$ at a point $p\in\mathcal{M}$ will be denoted by $\mathrm{I\!I}_{p}$ and is a symmetric bilinear map $\mathrm{I\!I}_{p}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\operatorname{N}_{p}\mathcal{M}$ intrinsic to the manifold $\mathcal{M}$. Fixing some $u\in\operatorname{N}_{p}\mathcal{M}$ we also define the directed second fundamental form $\mathrm{I\!I}_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\mathbb{R}:(v,w)\mapsto\left\langle\mathrm{I\!I}_{p}(v,w),u\right\rangle_{\operatorname{N}_{p}\mathcal{M}}$. The Weingarten map $S_{p}^{u}$ at a point $p\in\mathcal{M}$ in the direction $u\in\operatorname{N}_{p}\mathcal{M}$ is defined as the unique self-adjoint linear operator $S_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\to\operatorname{T}_{p}\mathcal{M}$ such that $\left\langle w,S_{p}^{u}(v)\right\rangle_{\operatorname{T}_{p}\mathcal{M}}=\mathrm{I\!I}_{p}^{u}(v,w)$ for all $v,w\in\operatorname{T}_{p}\mathcal{M}$. A useful operator that has been studied in Abatzoglou (1978); Breiding & Vannieuwenhoven (2021); Alvarez et al. (2004) is
@@
In Breiding & Vannieuwenhoven (2021) it has been shown (see Lemma 8) that $H_{x}$ is invertible on $\mathcal{T}$. Now we have the following useful identities that hold for $x\in\mathcal{T}$
@@ A.1 Properties of (16)
The operator (16) is invertible on $\mathcal{T}(\tau_{\mathcal{M}})$. Moreover, (16) satisfies [^1] In (Breiding & Vannieuwenhoven, 2021, Theorem 4.3) the quantity $\lVert H_{x}^{-1}\rVert$ has been shown to equal the condition number of a certain critical point problem associated with $\mathcal{M}$.
@@
In (Breiding & Vannieuwenhoven, 2021, Lemma A.2) the following condition has been established: Let $\mathcal{S}=\{(a,p)\in\mathbb{R}^{n}\times\mathcal{M}\mid a-p\in\operatorname{N}_{p}\mathcal{M}\}$. Then $\mathcal{S}$ is diffeomorphic to the normal bundle $\operatorname{N}\mathcal{M}$ via the diffeomorphism $\Phi:\operatorname{N}\mathcal{M}\to\mathcal{S}:(v,p)\mapsto(p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v),p)$. Consider the operator $\Pi:\mathcal{S}\to\mathbb{R}^{n}:(a,p)\mapsto a$ and the domain where its differential is invertible $\mathcal{W}=\{(a,p)\in\mathcal{S}\mid\Pi^{\prime}(a,p):\operatorname{T}_{(a,p)}\mathcal{S}\to\mathbb{R}^{n}\text{\ invertible}\}$. Then $H_{x}$ is invertible iff $(x,\pi(x))\in\mathcal{W}$. But $\Pi\circ\Phi:\operatorname{N}\mathcal{M}\mapsto\mathbb{R}^{n}:(v,p)\mapsto p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v)$ being a diffeomorphism (and thus having an inveritble differential) is precisely the condition in the definition of the tubular neighborhood $\mathcal{T}$. The formulas for $P_{0}(x)$ in local coordinates as well as $\lVert P_{0}(x)\rVert$ are given in (Abatzoglou, 1978, Theorem 4.1, Corollary 4.1). To see that $P_{0}(x)$ is bounded on $\mathcal{T}(\tau)$ for any $\tau\in(0,\tau_{\mathcal{M}})$ it sufficies to note that in the tubular neighborhood $\mathcal{T}=\mathcal{T}(\tau_{\mathcal{M}})$ we always have $\lVert\pi(x)-x\rVert<\rho(\pi(x),x-\pi(x))$ and that $\overline{\mathcal{T}(\tau)}$ is a compact subset thereof. The second inequality follows from $1/\rho(p,u)\leq\kappa_{\mathcal{M}}$. ∎
@@
A.2 Densities on manifolds
Given a manifold $\mathcal{M}$ as in Section A and any chart $\psi:\mathcal{V}\to\mathcal{U}\subseteq\mathcal{M}$, the volume measure $\operatorname{Vol}_{\mathcal{M}}$ is uniquely defined by
@@ A.2 Densities on manifolds
Appendix B More on the Stein score function
@@
Appendix C Variance-exploding diffusion models
In the variance-exploding scheme of score-based diffusion models Song et al. (2020); Tang & Zhao (2025), one considers the following stochastic differential equation on the finite interval $[0,T]$:

### v2

@@ 7 Conclusion and Future Work
Acknowledgments and Disclosure of Funding
We thank the reviewers for their constructive suggestions. This work was supported as a part of NCCR Automation, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 51NF40_225155). Riccardo De Santi is supported by the ETH AI Center through an ETH AI Center doctoral fellowship.
Appendix A Connections and comparisons to prior work
A.1 Connection to denoising autoencoders
In the following we show how our results connect to learning denoising auto-encoders. The following theorem is shown in Alain and Bengio (2014):
Theorem 8 (Theorem 1 in Alain and Bengio (2014)).
Let $\mu$ be a probability measure on $\mathbb{R}^{d}$ and consider for $\sigma>0$ the reconstruction problem
$\min_{r}L_{\operatorname{DAE}}(r)\quad\text{for}\quad L_{\operatorname{DAE}}(r)=\mathbb{E}_{x\sim\mu\,,\;\epsilon\sim\mathcal{N}(0,I)}\|x-r(x+\sigma\epsilon)\|^{2}\,.$
If $\mu$ is absolutely continuous with (Lebesgue) density $p$, the optimal reconstruction function $r_{\sigma}^{*}$ satisfies
$r_{\sigma}^{*}(x)=\frac{\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}p(x-\sigma\epsilon)(x-\sigma\epsilon)}{\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}p(x-\sigma\epsilon)}\overset{\sigma\to 0}{=}x+\sigma^{2}\nabla\log p(x)+o(\sigma^{2})$
at any $x\in\mathbb{R}^{d}$ with $p(x)>0$.
To see how this result relates to Theorem 1, note that function $r_{\sigma}^{*}$ here is precisely our “approximate projection” $\pi_{\sigma}$, since the first above expression corresponds to the posterior of observing $x$ under the convolution of the prior $p$ with $\mathcal{N}(0,\sigma^{2}I)$, i.e. in our notation
$r_{\sigma}^{*}(x)=\mathbb{E}\nu_{x,\sigma}=\pi_{\sigma}(x).$
Therefore, if $\mu$ has the Lebesgue-density $p$ (i.e. the data manifold would have full dimension $k=d$), then Alain and Bengio (2014) show that
$x+\sigma^{2}\nabla\log p_{\sigma}(x)\overset{\sigma\to 0}{=}x+\sigma^{2}\nabla\log p(x)+o(\sigma^{2}).$
In particular, if $p(x)>0$, then the last equation states that
$\lim_{\sigma\to 0}x+\sigma^{2}\nabla\log p_{\sigma}(x)=x.$
Theorem 1 on the other hand deals with the case $k<d$, i.e. when the manifold is truly lower-dimensional. In this case, $\mu$ has no density $p$ w.r.t. the Lebesgue measure on $\mathbb{R}^{d}$, and it holds that
$\lim_{\sigma\to 0}x+\sigma^{2}\nabla\log p_{\sigma}(x)=\pi(x),$
where $\pi$ denotes the projection operation onto the data manifold $\mathcal{M}$. While Alain and Bengio (2014) does consider manifold support in their discussion (see (Alain and Bengio, 2014, Section 3.4, Figure 5)), note that the population data measure in this case is not absolute continuous w.r.t. the Lebesgue measure on $\mathbb{R}^{d}$, and hence $\log p(x)$ has no meaning. Consequently, results of Alain and Bengio (2014), e.g. the expansion above regarding $r_{\sigma}^{*}$, are not applicable. Theorem 1 can therefore be seen a (uniform) generalization of Alain and Bengio (2014) to an order of $\widetilde{O}(\sigma)$, establishing the asymptotic convergence of $r_{\sigma}^{*}$ and its Jacobian for $\sigma\to 0$ as
$r_{\sigma}^{*}(x)=\pi(x)+\widetilde{O}(\sigma)\,,\quad(r_{\sigma}^{*})^{\prime}(x)=\pi^{\prime}(x)+\widetilde{O}(\sigma)\,.$
A.2 Trivializations and natural gradients using parametric manifold learning
In this section we compare our approach to latent space optimization. For this purpose, we first describe the latter problem in terms of trivializations of (1) using local parametrizations: Given a finite atlas $\mathcal{A}=(\psi_{i},\mathcal{U}_{i},\mathcal{V}_{i})_{i=1}^{m}$ of $\mathcal{M}$ with open $\mathcal{V}_{i}\subseteq\mathbb{R}^{k}$ and $\mathcal{U}_{i}\subseteq\mathcal{M}$ and charts $\psi_{i}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ it clearly holds
$\min_{x\in\mathcal{M}}f(x)=\min_{i=1,\ldots,m}\min_{z\in\mathcal{V}_{i}}f(\psi_{i}(z))\,.$
In the following we will focus on one inner subproblem by fixing a single chart $(\psi,\mathcal{U},\mathcal{V})$ and consider
$\min_{x\in\mathcal{U}}f(x)=\min_{z\in\mathcal{V}}f(\psi(z))\,.$ (16)
While for the above equality we only need $\psi$ to be surjective (i.e. $\psi(\mathcal{V})=\mathcal{U}$), for equivalence of first-order stationary points of both problems we need $\psi$ to be a submersion: To see this, note that stationarity of $x_{*}\in\mathcal{U}$ for the left hand side in equation 16 is equivalent to $\operatorname{grad}_{\mathcal{M}}f(x_{*})=\operatorname{P}_{\operatorname{T}_{x_{*}}\mathcal{M}}\nabla f(x_{*})=0$ and stationarity of any $z_{*}\in\mathcal{V}\cap\psi^{-1}(\{x_{*}\})$ for the right hand side is equivalent to $\psi^{\prime}(z_{*})^{\top}\nabla f(\psi(z_{*}))=\psi^{\prime}(z_{*})^{\top}\nabla f(x_{*})=0$. In particular, for these two notions to be equivalent a-priori for any $f$ and $x_{*}$, the columns of $\psi^{\prime}(z_{*})$ must span $\operatorname{T}_{x_{*}}\mathcal{M}$, which implies that $\psi$ is a submersion. In particular, if $\psi$ is a chart, then the stationary points of both sides in equation 16 are in a one-to-one correspondence and only then both problems are equivalent from an first-order optimization point of view.
Two possible ways to solve the right hand side in (16) are the latent space gradient flow (LSGF) and the pull-back Riemannian gradient flow (PBRGF) given by
$\dot{z}$ $=-\nabla(f\circ\psi)(z)=-\psi^{\prime}(z)^{\top}\nabla f(\psi(z))\,$, (17)
$\dot{z}$ $=-G_{\psi}(z)\psi^{\prime}(z)^{\top}\nabla f(\psi(z))\,$, (18)
respectively, where $G_{\psi}(z)=\psi^{\prime}(z)^{\top}\psi^{\prime}(z)$ is the metric tensor of $\psi$ at $z\in V$. Here PBRGF [^2] The Euler discretization of the continuous dynamics (18) is also known as natural gradient descent (Amari, 1998, Theorem 1) in the classical Riemannian optimization literature is obtained from LSGF via a change in inner product on the tangent space $\operatorname{T}_{\psi(z)}\mathcal{M}$ and is equivalent to the Riemannian gradient flow in the sense that for $x=\psi(z)$ it holds that
$\dot{x}=-\operatorname{grad}f(x)\,.$
While (17) and (18) are both possible approaches to solving (1) and have been applied in the past Pegios et al. ; Tripp et al. (2020), there are a few challenges present these methods:
  • Parameterizing the full atlas of a manifold is hard: The atlas $\mathcal{A}$ of a manifold typically consists of multiple charts $(\psi_{i},U_{i},V_{i})$. The number of charts as well as their span across the manifold is typically unknown and hard to estimate a-priori.
  • Intrinsic dimension of the manifold unknown: The dimension of the latent space $\mathcal{Z}$ must correspond to the manifold dimension, the estimation of which is not a trivial task.
  • Generative model training: The implementation of (18) or (17) require $\psi$ to be invertible and the inverse coordinate map $\psi^{-1}$ to be well approximated by a neural network. It is not straightforward to align this requirement with the standard training objectives commonly used in generative models, such as the likelihood-based loss in VAEs, the denoising score-matching loss in diffusion models, or the adversarial loss in GANs. Further approaches such as M-flows Brehmer and Cranmer (2020) require $\psi$ to be essentially a normalizing flow, which is fundamentally incompatible with common architectures such as U-Nets or transformers.
  • Expensive update rules for PBRGF: The update rule of PBRGF requires (i) computing the matrix $G_{\psi}(z)$, which in turn involves the explicit evaluation of the Jacobian $\psi^{\prime}(z)$, and (ii) inverting $G_{\psi}(z)$. Both operations become computationally expensive when $\psi$ is implemented by a neural network even for a moderate intrinsic dimension $k$.
  • Limited latent space validity region: Since $\mathcal{Z}$ is representing the local coordinates of a single chart, the validity of the model $\psi$ is only given in a bounded domain of $\mathcal{Z}$. Optimizing outside of this domain requires changing charts, since otherwise the obtained latent points will have no meaning in the data-space. Estimating the validity domain of $\psi$ is difficult a-priori Notin et al. (2021).
In the above light, the advantages of the algorithms DLF and DRGD described in Sections 4 and 5 are as follows:
  • Easy to parameterize with a neural network: The central object in our approach is the projection operator onto the data manifold $\mathcal{M}$, which is provably $C^{2}$ in the tubular neighborhood of any $C^{3}$ manifold. Assuming $\mathcal{M}$ is compact, standard universal approximation results ensure that the target map $\pi$ can be well-approximated by a single neural network
  • Dimension-agnostic algorithms: Both DLF and DRGD require no knowledge about the manifold dimension $k$.
  • Alignment with denoising score matching: DLF and DRGD use an approximation of $\pi$ via the score function of $p_{\sigma}$ (up to scaling and a residual term of $x$), so the denoising score-matching loss used to train diffusion models is directly aligned with our target quantity. Our approach imposes no architectural constraints: in our experiments, for example, we successfully use both MLPs and U-Nets.
  • Simple and efficient update rule The update rules (8) and (12) are highly compatible with modern hardware and software stacks and requires only a single forward–backward pass of the network per iteration (see Remark 4)
  • Convergence guarantees in the embedding space: Theorem 3 and 5 provide explicit convergence analysis with guarantees on both optimality and feasibility, stated directly for the original data manifold in the data space, where convergence is semantically meaningful.
A.3 Recovery of iterate complexity of Riemannian gradient descent
In this section we show that in the case of $\sigma=\epsilon=0$, Theorem 5 recovers the corresponding result for the Riemannian gradient descent. To see this, let us recall the standard complexity guarantee for Riemannian gradient descent under a general smooth non-convex objective $f$: In (Boumal, 2023, Corollary 4.9) it is stated that (taking $R_{x}(v)=\pi(x+v)$ as the retraction defined on some ball $B_{r}(x)\cap T_{x}\mathcal{M}$, where for us $r>0$ can be picked uniformly over $x\in\mathcal{M}$) if $f\circ\pi$ satisfies the pullback $L$-Lipschitz condition
$f(\pi(x+v))-f(x)\leq v^{\top}\nabla f(x)+\frac{L}{2}\|v\|^{2}\text{\ for all\ }v\in B_{r}(x)\cap T_{x}\mathcal{M}\,.$
then for step-size $\gamma_{k}=\frac{1}{L}$ the RGD generates iterates $(x_{k},v_{k}=-\gamma_{k}\pi^{\prime}(x_{k})\nabla f(x_{k}))$ that satisfy
$\min_{k=0,\ldots,N-1}\|\operatorname{grad}f(x_{k})\|\leq\frac{\sqrt{2L(f(x_{0})-f_{*})}}{\sqrt{N}}\,.$
provided that $v_{k}\in B_{r}(x)\cap T_{x}\mathcal{M}$ for all $k=0,\ldots,N-1$.
Now we set the following in Theorem 5 in our paper: $s=\pi$, $\sigma=0$. The pullback $L_{0}$-Lipschitz condition is satisfied under our assumption $L=\operatorname{Lip}(\nabla f)$ for a different constant $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi))$, which can be estimated in terms of $L$, $C$, $\tau$ and $M$ (see the proof of Theorem 5). The rate obtained in Theorem 5 reads then
$\min_{k=0,\ldots,N}\|\operatorname{grad}f(x_{k})\|^{2}\leq\frac{4D}{N\gamma_{\min}(1-L_{0}\gamma_{\max}/2)}\,$,
where $\gamma_{\min}(1-L_{0}\gamma_{\max}/2)$ can be lower-bounded by a constant constructed from $\tau$ and $L_{0}$. Moreover, $2D$ is simply an estimation of $f(x_{0})-f_{*}$ via the triangle inequality (see the proof of Theorem 5). Thus in this case both theorems guarantee an $O(1/\sqrt{N})$ best-norm-of-gradient convergence rate with similar constants.
Appendix B More on manifolds and distance functions
@@
In addition to the notation introduced in Section 2.1, we note that $\tau_{\mathcal{M}}$ is also the largest $\tau\geq 0$ such that the map $\{(p,v)\in\operatorname{N}\mathcal{M}\mid\lVert v\rVert<\tau\}\to\mathbb{R}^{d}:(p,v)\mapsto p+v$ is a diffeomorphism. By a tubular neighborhood of radius $\tau\in(0,\tau_{\mathcal{M}}]$ we mean a set of the form $\mathcal{T}(\tau)=\{p+v\mid p\in\mathcal{M}\,,v\in\operatorname{N}_{p}\mathcal{M}\,,\lVert v\rVert<\tau\}$. Moreover, for $x\in\mathbb{R}^{d}$ let $\operatorname{dist}_{\mathcal{M}}(x)=\inf_{p\in\mathcal{M}}\lVert x-p\rVert$ denote the distance function so that $\operatorname{d}(x)=\frac{1}{2}\operatorname{dist}_{\mathcal{M}}(x)^{2}$. The second fundamental form of $\mathcal{M}$ at a point $p\in\mathcal{M}$ will be denoted by $\mathrm{I\!I}_{p}$ and is a symmetric bilinear map $\mathrm{I\!I}_{p}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\operatorname{N}_{p}\mathcal{M}$ intrinsic to the manifold $\mathcal{M}$. Fixing some $u\in\operatorname{N}_{p}\mathcal{M}$ we also define the directed second fundamental form $\mathrm{I\!I}_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\mathbb{R}:(v,w)\mapsto\left\langle\mathrm{I\!I}_{p}(v,w),u\right\rangle_{\operatorname{N}_{p}\mathcal{M}}$. The Weingarten map $S_{p}^{u}$ at a point $p\in\mathcal{M}$ in the direction $u\in\operatorname{N}_{p}\mathcal{M}$ is defined as the unique self-adjoint linear operator $S_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\to\operatorname{T}_{p}\mathcal{M}$ such that $\left\langle w,S_{p}^{u}(v)\right\rangle_{\operatorname{T}_{p}\mathcal{M}}=\mathrm{I\!I}_{p}^{u}(v,w)$ for all $v,w\in\operatorname{T}_{p}\mathcal{M}$. A useful operator that has been studied in Abatzoglou (1978); Breiding and Vannieuwenhoven (2021); Alvarez et al. (2004) is
@@
In Breiding and Vannieuwenhoven (2021) it has been shown (see Lemma 10) that $H_{x}$ is invertible on $\mathcal{T}$. Now we have the following useful identities that hold for $x\in\mathcal{T}$
@@ B.1 Properties of (19)
The operator (19) is invertible on $\mathcal{T}(\tau_{\mathcal{M}})$. Moreover, (19) satisfies [^3] In (Breiding and Vannieuwenhoven, 2021, Theorem 4.3) the quantity $\lVert H_{x}^{-1}\rVert$ has been shown to equal the condition number of a certain critical point problem associated with $\mathcal{M}$.
@@
In (Breiding and Vannieuwenhoven, 2021, Lemma A.2) the following condition has been established: Let $\mathcal{S}=\{(a,p)\in\mathbb{R}^{n}\times\mathcal{M}\mid a-p\in\operatorname{N}_{p}\mathcal{M}\}$. Then $\mathcal{S}$ is diffeomorphic to the normal bundle $\operatorname{N}\mathcal{M}$ via the diffeomorphism $\Phi:\operatorname{N}\mathcal{M}\to\mathcal{S}:(v,p)\mapsto(p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v),p)$. Consider the operator $\Pi:\mathcal{S}\to\mathbb{R}^{n}:(a,p)\mapsto a$ and the domain where its differential is invertible $\mathcal{W}=\{(a,p)\in\mathcal{S}\mid\Pi^{\prime}(a,p):\operatorname{T}_{(a,p)}\mathcal{S}\to\mathbb{R}^{n}\text{\ invertible}\}$. Then $H_{x}$ is invertible iff $(x,\pi(x))\in\mathcal{W}$. But $\Pi\circ\Phi:\operatorname{N}\mathcal{M}\mapsto\mathbb{R}^{n}:(v,p)\mapsto p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v)$ being a diffeomorphism (and thus having an inveritble differential) is precisely the condition in the definition of the tubular neighborhood $\mathcal{T}$. The formulas for $P_{0}(x)$ in local coordinates as well as $\lVert P_{0}(x)\rVert$ are given in (Abatzoglou, 1978, Theorem 4.1, Corollary 4.1). To see that $P_{0}(x)$ is bounded on $\mathcal{T}(\tau)$ for any $\tau\in(0,\tau_{\mathcal{M}})$ it sufficies to note that in the tubular neighborhood $\mathcal{T}=\mathcal{T}(\tau_{\mathcal{M}})$ we always have $\lVert\pi(x)-x\rVert<\rho(\pi(x),x-\pi(x))$ and that $\overline{\mathcal{T}(\tau)}$ is a compact subset thereof. The second inequality follows from $1/\rho(p,u)\leq\kappa_{\mathcal{M}}$. ∎
@@
B.2 Densities on manifolds
Given a manifold $\mathcal{M}$ as in Section B and any chart $\psi:\mathcal{V}\to\mathcal{U}\subseteq\mathcal{M}$, the volume measure $\operatorname{Vol}_{\mathcal{M}}$ is uniquely defined by
@@ B.2 Densities on manifolds
Appendix C More on the Stein score function
@@
Appendix D Variance-exploding diffusion models
In the variance-exploding scheme of score-based diffusion models Song et al. (2020); Tang and Zhao (2025), one considers the following stochastic differential equation on the finite interval $[0,T]$: